OpenStudy (anonymous):

There are 8 electrical engineers and 12 computer engineers who are doing a group project. We need to divide them into 4 groups of 5 so that each group contains exactly 2 electrical engineers and 3 computer engineers. How many ways can this be done?

OpenStudy (anonymous):

You want to split up the problem into simple, easily counted tasks.

OpenStudy (anonymous):

I'd first assign each group their electrical engineers, then assign them their computer engineers.

OpenStudy (anonymous):

so 8! / 6! + 12! / 9!

OpenStudy (anonymous):

These are sequential tasks, so you multiply their counts...

OpenStudy (anonymous):

8 ways

OpenStudy (anonymous):

i don't understand how you got to 8 ways.

OpenStudy (anonymous):

495 is the answer I just checked on the calculator

OpenStudy (anonymous):

Yes, but how exactly do you get that?

OpenStudy (anonymous):

For assigning electrical engineers, I would try: \[ \binom{8}{2}\cdot \binom{6}{2}\cdot \binom{4}{2}\cdot \binom{2}{2} \]

OpenStudy (anonymous):

For assigning computer engineers, I would try: \[ \binom{12}{3}\cdot \binom{9}{3}\cdot \binom{6}{3}\cdot \binom{3}{3} \]

OpenStudy (anonymous):

The total count would be the product of these two.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

Most of it. Why would you multiply rather than add?

OpenStudy (anonymous):

You add when you have an option of doing either task. You multiply when you have two do both tasks.

OpenStudy (anonymous):

That would make sense since in this question it's an "and".